2. What applied horizontal force is required to accelerate a 5 kg box along a horizontal surface with an acceleration of 2 m/s2 if the coefficient of friction is 0.15?


1
Expert's answer
2021-09-20T10:00:15-0400

Given:

"m=5\:\rm kg"

"a=2\:\rm m/s^2"

"\mu=0.15"


The Newton's second law says

"ma=F-F_f"

"ma=F-\mu mg"

Hence, the requared force

"F=m(a+\mu g)=5*(2+0.15*9.8)=17.4\:\rm N"